Please help with me
Find the fractions equal to the given decimals
1. 0.55555

Answer by jim_thompson5910(35256)

x = 0.55555
10x = 5.5555


10x - x = 5.5555 - 0.55555

9x = 5

x = 5/9

Answer: 5/9 = 0.555555....